# Galley: Modern Query Optimization for Sparse Tensor Programs

### v4

@@ 7.3.2. Degree Statistics and the Chain Bound
Galley stores degree statistics as the default, and it uses them to compute upper bounds on the number of non-fill entries in intermediate expressions. A degree statistic, denoted as $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, given a matrix $A_{ij}$, then $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding that has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
Constructor
This function first computes the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, it sums over the $X$ dimensions and takes the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$.
Map (Annihilating)
Annihilating map operations serve as conjunctive queries with respect to the sparsity patterns of the inputs. Therefore, any degree statistics that are valid for an input are also valid about the output. This function computes the statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by a union
$\mathcal{D}_{C}=\bigcup_{j}A^{(j)}_{I_{j}}$
Map (Non-Annihilating)
In this case, the function must be especially careful to ensure that it maintains the upper bounds. First, it extends the degree constraints from each input to cover the full set of indices. For example, if we have $DC_{X}(i|j)$ and want to extend it to the dimension $k$, the function computes $D_{X}(ik|j)=DC(i|j)*n_{k}$. Then, it computes degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition:
@@
Estimator
This function calculates an upper bound (eq. performs sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Its search begins with the empty set; it then uses a breadth-first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bounds the number of non-zeros in the result.
@@ 8. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) ML algorithms over joins, (2) subgraph counting, and (3) breadth-first search. We choose those workloads because they exercise different aspects of our optimizer on real-world use-cases: ML algorithms over joins require careful logical optimizations over programs with mixtures of dense and sparse inputs and non-linear operators; subgraph counting requires both logical and physical optimization of complex sum-product expressions over highly sparse inputs and demonstrates Galley’s advantage over a relational engine even for very sparse workloads; breadth-first search requires careful selection of tensor formats over the course of the computation, showing the benefit of physical optimization for even simple computations. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while requiring only a concise, declarative input program from the user. Overall, we show that Galley:
[…]
@@
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://anonymous.4open.science/r/Galley-21BF/. We used the sparse tensor compiler Finch [^2] https://github.com/FinchTensor/Finch.jl for execution, and all experiments are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 7.3.2 for sparsity estimation. Experiments for all methods are run five times, and the mean execution time is reported. We perform all experiments on a warm cache, and we separately report the compilation and optimization times.
8.1. Machine Learning Algorithms
To explore end-to-end program optimization, we experiment with simple ML algorithms over joins, represented entirely in tensor algebra. For this, we use the TPC-H benchmark at a scale factor of .25.25.25.25 in order to keep the comparison methods in memory [^3] Galley’s execution tends to be less memory-intensive, and we have successfully run it at scale factor 1.5.. We consider two join queries on this dataset: star and self-join. First, we perform a star join over the line items table to gather features about suppliers, parts, orders, and customers. This is expressed as follows, where $L,S,P,O,$ and $C$ are tensors representing the line items, suppliers, parts, orders, and customers tables, respectively:
@@ 8.1. Machine Learning Algorithms
Given these definitions, we consider a range of ML algorithms: (1) linear regression inference, (2) logistic regression inference, (3) covariance matrix calculation, and (4) neural network inference. For comparison, we also implement two versions of each of these using the Finch compiler. The dense version uses a fully dense format to represent the features, whereas the sparse version uses a sparse level for the features to compress the one-hot encoding. We also implemented these algorithms using the Pandas and Numpy libraries for reference.
@@
These algorithms stress the ability of Galley to handle both combinations of complex operators and combinations of sparse and dense inputs. The definitions of the feature tensors combine pointwise multiplication and addition, and algorithms like logistic regression and neural networks wrap these definitions in more non-linear operators (e.g. relu and sigmoid) and aggregates. Further, while the line item tensor is highly sparse, both the feature and parameter tensors are moderately to fully dense. In all of these examples, Galley is able to manipulate these complex expressions to avoid materializing large, sparse intermediates. Instead, its logical optimizer aggressively pushes the computation and aggregation down to the smaller, denser inputs. Consider logistic regression over the star query feature matrix. In this case, Galley distributes the parameter vector over the addition and performs the summations $\sum_{j}S_{sj}\theta_{j},\sum_{j}P_{pj}\theta_{j},\sum_{j}O_{oj}\theta_{j},$ and $\sum_{j}C_{cj}\theta_{j}$ before handling the remaining operation in one kernel. By doing this, it materializes a dense vector at every point and removes the feature dimension before operating on the large, sparse line item tensor.
@@
Execution Time. Fig. 6 shows that the execution time of Galley’s optimized programs is .5300×.5-300$\times$.5 - 300 × faster than the sparse Finch implementation. For the regression/neural network problems, this stems from pushing the parameter vector/matrix down to the feature matrices as discussed above. For the covariance calculation over the self join, Galley fully distributes the multiplication over the addition. Each distributed term is the product of four copies of the lineitem table $l_{i_{1},s_{1},p_{1}},l_{i_{2},s_{2},p_{1}}l_{i_{1},s_{3},p_{2}}l_{i_{2},s_{4},p_{2}}$ and one feature matrix from each $X$, e.g. $S_{s_{1},j}P_{p_{1},k}$. Within each term, Galley performs the summation over $i_{1}$ and $i_{2}$ first, before handling the summation of $p_{1}$ or $p_{2}$. Because $i_{1}$ and $i_{2}$ are keys, unlike $p_{1}$ and $p_{2}$, these initial sums produce small, dense intermediates, speeding up the whole computation. Lastly, the range of Galley’s execution times for the five iterations of each experiment was never more than $25\%$ of the mean.
@@
Optimization Time. Fig.  6 also shows that Galley’s optimizer has a reasonable overhead in this setting. For all experiments except for Covariance (SQ), Galley is faster than competing methods even when including the optimization time. Concretely, optimization takes .$15-3.0$ seconds on these workloads. The covariance calculations take longer to optimize due to the number of terms produced when they are fully distributed.
Figure 7. Subgraph Counting Execution Time
Figure 8. Subgraph Counting Optimization Time
Figure 9. Subgraph Counting Compilation Time
Figure 10. Sparsity Estimator Comparison
@@ 8.2. Subgraph Counting
In this section, we test Galley’s ability to optimize programs with a large number of highly sparse inputs. To do this, we implement a few common sub-graph counting benchmarks. The conversion from sub-graph counting to sparse tensor algebra is straightforward and results in traditional einsum expressions. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$; we can represent the count as
@@
To handle vertex labels, we add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study”(Sun and Luo, 2020; Park et al., 2020). We restrict the latter benchmark to query graphs with up to 8 vertices (hence the ”_lite” suffix). Because this is a relational workload, we compare it with DuckDB, a modern OLAP database (Raasveldt and Mühleisen, 2019). To separately discern the impact of logical vs physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. Finally, we provide results for the greedy logical optimizer, as well.
@@
Logical Optimization. Fig. 7 shows the execution time for each method and benchmark. The comparison between ‘DuckDB‘ and ‘Galley + DuckDB Backend‘ demonstrates the benefits of Galley’s logical optimizer. Galley’s logical optimizer breaks down the program into a series of aggregations which effectively minimize the necessary computation and materialization, resulting in a dramatic performance improvement. This has the largest impact on graphs with high degree values and/or high skew like the social network graphs, ‘dblp_lite‘ and ‘youtube_lite‘. In these cases, not pushing down aggregates can result in very large intermediate results. On low-degree graphs like ‘aids‘, these optimizations have less impact. Further, DuckDB hits the 300 second timeout on 50 out of 120 queries in the youtube_lite benchmark. In contrast, Galley never times out across all workloads.
@@
Physical Optimization. The impact of Galley’s physical optimizer can be seen by comparing ‘Galley‘ with ‘Galley + DuckDB Backend‘. By producing efficient STC kernels, Galley’s median execution is up to 8x faster than DuckDB even with the same logical plan. This shows that Galley is selecting efficient loop orders and formats which allows it to take advantage of the highly flexible language of STCs.
@@
Optimization Time. Fig. 8 shows the mean optimization time for each method on each workload. Galley has a mean optimization time of less than .15.15.15.15 seconds across all workloads, approaching the time taken by the highly efficient DuckDB optimizer.
@@
Compilation Time. Because it performs compilation using Finch at runtime, Galley incurs a compilation latency the first time that it invokes each unique STC kernel. Fortunately, these kernels are automatically cached by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 9. On the simpler workloads, which often reuse kernels, this cost is minimal. However, the more complex workloads both reuse kernels less and require compiling more complex kernels, significantly increasing compilation time. This suggests a need for faster STC compilation or the use of an interpreted engine for complex queries.
Sparsity Estimation. Finally, in Fig. 10, we use the sub-graph matching workloads to compare sparsity estimators and their effect on performance. Across all workloads, we see that the chain bound significantly out-performs the uniform estimator. This is because graph datasets tend to be skewed and correlated. Vertices have drastically different edge counts, and vertices with many edges tend to connect to vertices with many edges. This violates the implicit assumptions of the uniform estimator, leading it to optimistically materialize large intermediates. Due to this, the uniform estimator runs out of memory on many queries in the ‘youtube_lite‘ workload, so it is not pictured in this figure.
@@ 8.3. Breadth-First Search
To demonstrate the importance of selecting optimal physical formats for tensors, we implement a simple breadth-first search algorithm using Galley and different hand-coded Finch implementations. Both systems receive a single iteration at a time, and the total execution time across all iterations is reported. The core computation for each iteration is a masked sparse matrix times sparse vector operation that computes the new frontier vector. The main optimization decision is the visited and frontier vectors’ formats. The former grows monotonically over the course of the algorithm, while the number of non-zeros in the latter forms a curve, peaking in the middle iterations. We provide two implementations of Finch, using either a sparse or a dense vector for both intermediates. Fig. 11 shows that Galley is significantly faster than both dense and sparse approaches for 4 of the 5 graphs and is competitive for all graphs. During a single run of BFS, Galley begins with a sparse visited vector before switching to bytemap and dense formats as more nodes are discovered. The frontier vector begins as sparse, becomes dense, then again becomes sparse for the last few iterations as most nodes are deactivated. For 4/5 graphs, the total optimization time for Galley (not depicted in the figure) is less than .25.25.25.25 seconds. RoadNet has a very wide diameter leading to many iterations and a total optimization time of $4$ seconds. This experiment demonstrates the utility of sparsity-aware format selection, and future work should consider ways to amortize optimization time for iterative workloads.
9. Conclusion and Limitations
This paper introduced Galley, a system for efficient declarative sparse tensor programming. We describe and then demonstrate how it optimizes high-level program structure with its logical optimizer and how it lowers that program to an efficient implementation with its physical optimizer. Guiding these decisions are sparsity estimates of intermediate expressions, and we show that these estimates can be computed for arbitrary tensor algebra programs by implementing a minimal 5-function interface. Finally, we present results from our evaluation of experiments with Galley on ML algorithms over structured feature data, sub-graph counting, and breadth-first search.
We are excited to enrich Galley further with new optimizations in the future. Currently, Galley lacks support for complex loop structures (e.g., a single outer FOR loop that wraps multiple inner FOR loops) and parallelism. Both areas could benefit from cost-based optimization. Similarly, Galley does not consider hard memory constraints during optimization, but our use of cardinality bound methods provides an avenue for addressing this in future work.
Acknowledgements.
This work was supported by NSF SHF 2312195, NSF IIS 2314527, and DARPA SBIR: HR001123C0139.

### v5

@@ 6.3.2. Degree Statistics and the Chain Bound
Galley stores degree statistics by default uses them to compute upper bounds on tensors’ sparsities. A degree statistic, denoted as $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, given a matrix $A_{ij}$, $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding that has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
Constructor. This function first computes the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, it sums over the $X$ dimensions and takes the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$.
Map (Annihilating). Annihilating map operations can only reduce the degree for any $X$, $Y$ pair. Therefore, every input’s degree statistics are also valid for the output. If the inputs are $A^{(}1)_{I_{1}},...,A^{(}k)_{I_{k}}$, then the output’s statistics are,
$\mathcal{D}_{C}=\bigcup_{j}\mathcal{D}_{A^{(j)}_{I_{j}}}$
Map (Non-Annihilating). In this case, Galley extends the degree constraints from each input to cover the full set of indices. Then, it computes degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition:
@@
Estimator. This function calculates an upper bound (eq. performs sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Its search begins with the empty set; it then uses a breadth-first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bounds the number of non-zeros in the result.
Example 0.
Suppose $A_{ij}$ and $B_{jk}$ are $100x100$ sparse matrices with $D_{A}(ij|\emptyset)=1000,D_{A}(i|j)=10,D_{B}(jk|\emptyset)=200$, and we want to bound $nnz(\sum_{j}A_{ij}B_{jk})$. Because multiplication is an annihilating operation in this case, the degree constraints of $\sum_{j}A_{ij}B_{jk}$ are simply the union of the constraints for $A$ and $B$. To get a bound, we start by conditioning on the empty set and try to reach the output’s index set, $i,k$, via the constraints, e.g. $D_{B}(jk|\emptyset)*D_{A}(i|j)=2000$.
@@
(a) Matrix Chain Multiplication (ABC)
(b) Elementwise Matrix Multiplication (A*B*C)
(c) Sum of Matrix Chain (SUM(ABC))
Figure 7. Linear Algebra Kernels
Table 1. Experimental Dataset Sizes
Dataset Size
TPCH (SF .25 - SF 5.0) .3-6 GB
aids 11 MB
human 1.5 MB
yeast 1.2 MB
dblp 21 MB
youtube 63 MB
Epinions 5.1 MB
Kron 34 MB
LiveJournal .5 GB
Orkut 1.7 GB
RoadNet 41 MB
@@ 7. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) ML algorithms over joins, (2) unstructured sparse linear algebra (3) subgraph counting, and (4) breadth-first search. We choose those workloads because they exercise different aspects of our optimizer on real-world use-cases: ML algorithms over joins require careful logical optimizations over programs with mixtures of dense and sparse inputs and non-linear operators; core linear algebra expressions demonstrate the broad utility of Galley; subgraph counting requires both logical and physical optimization of complex sum-product expressions over highly sparse inputs and demonstrates Galley’s advantage over a relational engine even for very sparse workloads; breadth-first search requires careful selection of tensor formats over the course of the computation, showing the benefit of physical optimization for even simple computations. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while requiring only a concise, declarative input program from the user. Overall, we show that Galley:
  • Performs logical optimizations resulting in $\textbf{1-300}\times$ faster execution for ML algorithms over joins compared to hand-optimized and Pandas implementations and $\textbf{.5-20}\times$ faster runtime when including optimization.
  • Optimizes in a mean time of at most 0.1 seconds for all subgraph counting workloads, with $\textbf{5-20}\times$ faster median execution than DuckDB.
@@
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://github.com/kylebd99/Galley.jl. We used the sparse tensor compiler Finch [^2] https://github.com/FinchTensor/Finch.jl for execution, and all experiments are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 6.3.2 for sparsity estimation. Experiments for all methods are run five times, and the mean execution time is reported. We perform all experiments on a warm cache, and we separately report the compilation and optimization times.
7.1. Machine Learning Over Joins
To explore end-to-end program optimization, we experiment with simple ML algorithms over joins. This represents a typical machine learning use case where the feature matrix is constructed from a variety of tables stored within an enterprise database that are joined together before training. Prior work has shown that co-optimizing these mixed RA/LA problems can yield significant benefits (Schleich et al., 2016; Kumar et al., 2015; Chen et al., 2016; Park et al., 2022; Grigorov et al., 2023). For this, we consider two join queries over the TPC-H benchmark: a star join and a self join, at scale factor 5 and .25, respectively. The star join is expressed as follows, where $L,S,P,O,$ and $C$ are tensors representing the line items, suppliers, parts, orders, and customers tables, respectively:
@@ 7.1. Machine Learning Over Joins
We consider a range of ML algorithms: (1) linear regression inference, (2) logistic regression inference, (3) covariance matrix calculation, and (4) neural network inference. We implement two versions of each of these using the Finch compiler. The dense version uses a dense feature matrix, and the sparse version uses CSR matrix to compress the one-hot encoding. We also implemented two standard baselines; 1) using Pandas for the joins and Numpy for the linear algebra 2) using Polars for the joins and PyTorch for the linear algebra. The latter supports parallelism, so we’ve included parallel results for it as well (marking the parallel bars with P and serial bars with S).
@@
These algorithms stress the ability of Galley to handle complex operators and combinations of sparse and dense inputs. The definitions of the feature tensors combine pointwise multiplication and addition, and algorithms like logistic regression and neural networks wrap these definitions in non-linear operators (e.g. relu and sigmoid) and aggregates. Further, while the line item tensor is highly sparse, both the feature and parameter tensors are moderately to fully dense.
@@
Execution Time. Fig. 6 shows that the execution time of Galley’s optimized programs is $5-300\times$ faster than the sparse Finch implementation. For the regression/neural network problems, this stems from pushing the multiplication with the parameter vector/matrix down to the feature matrices. For the covariance calculation over the self join, Galley fully distributes the multiplication over the addition then aggregates away the sparse $i_{1},i_{2}$ dimensions. This produces small, dense intermediates which can be used to calculate the covariance efficiently.
@@
Optimization Time. Fig.  6 also shows that Galley’s optimizer has a reasonable overhead in this setting. Concretely, optimization takes .55.0.5-5.0 seconds across all workloads.
Table 2. Total Subgraph Counting Execution Time (S)
Workload Galley (Greedy) Galley+DuckDB DuckDB Umbra 1 (24)
human .17 (.43) .156 .12 .04 (.02)
aids 32.1 (29.43) 43.32 78.16 7.47 (1.89)
yeast_lite 2.96 (3.85) 8.91 1633 367.28 (51.56)
dblp_lite 32.31 (30.44) 39.31 3294 75.51 (18.23)
youtube_lite 240.24 (219.47) 1591.27 17203 14208 (13866)
(a) Subgraph Counting Execution Time
(b) Subgraph Counting Optimization Time
(c) Sparsity Estimator Comparison
(d) Subgraph Counting Compilation Time
Figure 8. Subgraph Counting Experiments
7.2. Linear Algebra Kernels
In Fig. 7, we show how Galley can provide significant benefits even for simple workloads without structured data. In these experiments, $A$, $B$, and $C$ are $2000\times 2000$ uniformly sparse matrices. $A$ and $B$ have a density of .1.1, and the density of $C$ varies on the X axis. In the first experiment, Fig. 7(a), Galley improves on PyTorch’s execution in two ways: 1) when the matrices are less sparse, it chooses fully dense formats to store then intermediates and outputs 2) when $C$ is heavily sparse, it uses a right-to-left execution strategy for the chain, i.e. $(A(BC))$. In the second experiment, Fig. 7(b), Galley is able to fuse the computation when PyTorch is unable to, removing an intermediate materialization. Further, when $C$ is highly sparse, Galley iterates over the non-zeros of $C$ and looks them up in $A$ and $B$, rather than doing a symmetric intersection algorithm. In the third experiment, Fig. 7(c), Galley’s logical optimizer pushes down the outer summation to $A$ and $C$ first which avoids doing any expensive matrix multiplications. In almost all cases, Galley’s optimizations even overcome the benefits of parallelism when PyTorch is provided with 24 cores.
@@ 7.3. Subgraph Counting
In this section, we stress test Galley’s ability to optimize programs with a large number of highly sparse inputs by implementing several sub-graph counting benchmarks. These workloads represent the far end of the complexity and sparsity spectrum for sparse tensor compilers. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$; we can represent the count as,
@@
We add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study” (Sun and Luo, 2020; Park et al., 2020). We restrict them to query graphs with up to 8 vertices and 24 edges. Because this is a relational workload, we compare it with DuckDB and Umbra, two state-of-the-art modern OLAP databases (Raasveldt and Mühleisen, 2019; Neumann and Freitag, 2020). The latter is known to be one of the fastest databases for complex joins and aggregations, and we include both serial and parallel execution (Cli, [n. d.]). We did not include these systems in our other experiments due to the difficulty of framing the problems in SQL and because prior work has already demonstrated their challenges on pure LA workloads (Schüle et al., 2023). To separately discern the impact of logical vs physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. We also provide results for the greedy logical optimizer.
@@
Logical Optimization. Fig. 8(a) shows the execution time for each method and benchmark. The comparison between ‘DuckDB‘ and ‘Galley + DuckDB Backend‘ demonstrates the benefits of Galley’s logical optimizer. Galley’s logical optimizer breaks down the program into a series of aggregations which minimizes the necessary computation and materialization. This has the largest impact on graphs with high skew like the social network graphs, ‘dblp_lite‘ and ‘youtube_lite‘. In these cases, pushing down aggregates avoids very large intermediate results. DuckDB hits the 300 second timeout on 56 out of 120 queries in the youtube_lite benchmark, as does Umbra on 46 queries. In contrast, Galley never times out across all workloads.
@@
Physical Optimization. The impact of Galley’s physical optimizer can be seen by comparing ‘Galley‘ with ‘Galley + DuckDB‘. Galley’s median execution is up to 8x faster than DuckDB even with the same logical plan. This shows that Galley is selecting efficient loop orders and formats, effectively leveraging STCs.
@@
Optimization Time. Fig. 8(b) shows the mean optimization time for each method on each workload. Galley has a mean optimization time of less than .15.15 seconds across all workloads, faster than Umbra’s optimizer for 2 workloads.
@@
Compilation Time. Because it performs compilation at runtime, Galley incurs a compilation overhead when it invokes an STC kernel. These kernels are cached by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 8(d). On the simpler workloads, which often reuse kernels, this cost is lower. More complex workloads reuse kernels less, significantly increasing compilation time.
Comparing Figures 8 and 9, Galley’s optimization overhead is minimal (generally less than 1%) compared to the compilation overhead. Reducing this requires performance improvements to the underlying compiler which are out of scope for this work. Fortunately, the Finch project is working to improve this in two ways (1) by caching compilation to disk and (2) by migrating to the MLIR compiler infrastructure. As these improvements are made, Galley will immediately reap the benefits.
Sparsity Estimation. Finally, in Fig. 8(c), we use the sub-graph matching workloads to compare sparsity estimators and their effect on performance. Across all workloads, we see that the chain bound has significantly better tail performance. This is because it encourages more conservative query plans which better handle correlated and skewed queries/datasets.
@@ 7.4. Breadth-First Search
To demonstrate the importance of format selection, we implement a breadth-first search algorithm using Galley and hand-coded Finch implementations. Both systems receive a single iteration at a time, and the total execution time is reported. The core computation is a masked sparse matrix times sparse vector to compute the new frontier vector. The main decision is the visited and frontier vectors’ formats. The former’s sparsity grows monotonically over iterations, while the latter peaks in the middle iterations. We provide two implementations of Finch, using either a sparse or a dense vector for both. Fig. 9 shows that Galley’s mixture of sparse and dense formats is significantly fastest for 4 of the 5 graphs and is competitive for all graphs. For 4/5 graphs, the total optimization time (not depicted in the figure) is less than .25.25 seconds. This experiment demonstrates the utility of sparsity-aware format selection, and future work should consider ways to amortize optimization time for iterative workloads.
8. Related Work
9. Limitations
We are excited to enrich Galley with new optimizations in the future. Currently, Galley lacks support for complex loop structures (e.g., a single outer FOR loop that wraps multiple inner FOR loops), higher order functions (e.g. matrix inversion) or parallelism. However, we believe that these areas could benefit from cost-based optimization. Similarly, Galley does not consider hard memory constraints during optimization, but our use of cardinality bound methods provides an avenue for addressing this in future work.